OpenStudy (anonymous):

cosa-sina+1/cosa+sina-1=coseca+cota

OpenStudy (anonymous):

\[\frac{( 1+\cos a)-\sin a }{ \sin a-\left( 1-\cos a \right) }\] \[\frac{ 2\cos ^{2}\frac{ a }{ 2 }-2\sin \frac{ a }{ 2 }\cos \frac{ a }{2 } }{2\sin \frac{ a }{ 2 }\cos \frac{ a }{ 2 }-2\sin ^{2 }\frac{ a }{2 } }\]

OpenStudy (anonymous):

\[=\frac{ 2\cos \frac{ a }{ 2}\left( \cos \frac{ a }{ 2 } -\sin \frac{ a }{ 2 }\right) }{ 2\sin \frac{ a }{ 2 }\left( \cos \frac{ a }{ 2 }-\sin \frac{ a }{ 2 } \right) }=\cot \frac{ a }{ 2 }\]